# Structured Learning of Compositional Sequential Interventions

### v1

@@ Appendix B Proofs
$=Q_{C}(1-\alpha-{\rm gap})$ (38)
where ${\rm gap}=\frac{\sqrt{N}}{Z}\sqrt{\sum_{n=1}^{N}\left(w_{n}^{t^{\prime}}-{\rm E}(w^{t^{\prime}})\right)^{2}}=\frac{N}{Z}{\rm Var}(w^{t^{\prime}})=\frac{{\rm Var}(w^{t^{\prime}})}{{\rm E}(w^{t^{\prime}})}$, with ${\rm E}(w^{t^{\prime}})$ and ${\rm Var}(w^{t^{\prime}})$ being the empirical mean and variance of $w^{t}$ estimated on the $N$ calibration samples. Under the theorem assumptions, we have $w_{n}^{t^{\prime}}=(1-\epsilon)+\epsilon\frac{\delta(S_{n}^{t^{\prime}})}{p_{C}(S_{n}^{t^{\prime}})}$. Then, ${\rm E}(w^{t^{\prime}})=1-\epsilon+\epsilon\frac{{\rm E}(\delta(S^{t^{\prime}}))}{p_{C}(S^{t^{\prime}})}\geq 1-\epsilon$ and ${\rm Var}^{2}(w^{t^{\prime}})={\rm E}(w^{t^{\prime}}-{\rm E}(w^{t^{\prime}}))^{2}\leq\epsilon^{2}{\rm Var}^{2}(\frac{\delta(S^{t^{\prime}})}{p_{C}(S^{t^{\prime}})})\leq\frac{\epsilon^{2}}{(\min_{S}p_{C}(S^{t^{\prime}}))^{2}}{\rm Var}^{2}(\delta(S^{t^{\prime}}))\leq\frac{\epsilon^{2}}{(\min_{S}p_{C}(S^{t^{\prime}}))^{2}}$. The statement follows from
${\rm Prob}(Z\leq a)\geq{\rm Prob}(Z\leq b)\quad{\rm if}\quad a\geq b$ (39)
@@
$\alpha_{*}=\alpha+{\rm gap}\leq\alpha+\frac{\sqrt{{\rm Var}^{2}(w^{t^{\prime}})}}{{\rm E}(w^{t^{\prime}})}\leq\frac{\epsilon}{\min_{S}p_{C}(S)}\frac{1}{1-\epsilon}.$ (40)
@@ Appendix C Simulators
We design two different simulators: (1) a fully-synthetic simulator; and (2) a semi-synthetic simulator.
@@ C.1 General Setting
The simulators serve as fully controllable oracles to allow us test the performance of causal inference problems. In particular, we have the following parameters:
@@
  • $r$: the size of latent variable dimensions.
  • $z_{\text{dim}}$: the size of time invariant feature dimensions.
@@
Whenever possible, we set the same random seeds of $1,2,3,4,5$ to aid reproducible of our results. For the fully-synthetic simulator, a different seed indicates that it is a different simulator (random drawn of simulator parameters based on this seed) and also reflect the randomness comes from neural network parameter initialization and data splitting when training models. For the semi-synthetic simulator, the simulator parameters $\phi$ and $\beta$ are learned from real world data, but we will need to also draw synthetic parameters for $\psi$. In both cases, using different random seeds can be considered to drawing new problems where each has unique parameters.
@@ C.2 Fully-Synthetic Simulator
The first simulator is purely synthetic where all parameters are randomly sampled from some pre-defined distribution. The overall data generation process exactly follows the model we specified in Section 2.
We aims to generate a synthetic dataset for training with the following parameters: $r=5$, $T=20$, $T_{0}=10$, $K=5$, $N=50000$, and $z_{\text{dim}}=5$. We generate $200000$ at first place, but only use the first $50000$ samples for training. For testing our performance, we generate additional non-overlapping $M=5000$ samples with the same parameters, but this time with $T=25$. We make predictions on the last $5$ time steps based on the first $20$.
@@
$w_{3d}\sim\mathcal{U}(2,3).$
For $d=3,4$:
@@
We set $w_{1d}=0$, $w_{2d}=0$ and $w_{3d}=1$ to reflect the “idle” or “default” intervention level $d=0$. The values are designed to symmetric along the default intervention value. We generate random time series $\{D_{n}^{t}\}$ based on the identification results in Section 2.2, with the following rules:
@@
    for each unique unit, the same type of intervention action does not repeat (by sampling without replacement). We additionally assume at least $I>K/2$ (in this case, $I=3$) of the interventions show up in each unit-level time series.
@@
    once a $d>0$ intervention happens, there must be at least two consecutive $d=0$ interventions (see Theorem 1).
@@
    in the test dataset, at prediction time, we assume an unseen interventions in the first $20$ time series show up at exactly $21$, and then the time series has no intervention until the end. An example of this will be training $D=[0,0,\cdots,1,0,\cdots,0,2,0]$ of $T=20$ and testing $D=[0,0,\cdots,4,0,\cdots,0,3,0]+[2,0,0,0,0]$ of $T=25$.
We generate a time series $\{\psi_{n}^{t}\}$, based on Eq. 2, using the corresponding $w_{1d},w_{2d},w_{3d}$ values. We generate $\beta_{n}\sim\mathcal{N}(\mu_{n},\sigma_{n})$ for each user, each with a size of $N\times r$, as follows:
@@
We generate unit-specific covariates $\{Z_{n}\}$ with $z_{\text{dim}}=5$ of size $N\times z_{\text{dim}}$, as:
@@
This is then used to set up an initial $X_{n}^{0}$ (the initial starting point of the time series) with a random-parameter MLP with input dimension of $z_{\text{dim}}$ and output dimension of $1$. Finally, to create the synthetic time series $\{X_{n}^{0:T}\}$ (00 comes from the initial starting point), we randomly generate a neural network consisting of a single layer long-short term memory (LSTM) and a MLP to generate $\phi$ at each time step. We further compose $\phi$ with a sigmoid function so that it is bounded between $[-1,1]$. We iteratively sample the $\{X_{n}^{t}\}$ points with the conditional mean from Eq 1 (based on $X_{n}^{0:t-1}$, $Z_{n}$, and $D_{n}^{1:t}$) plus a standard Gaussian additive error term at each step.
The outcome of this simulator gives us the following data: time series $\{X_{n}^{0:T}\}$, intervention sequence $\{D_{n}^{1:T}\}$ and covariates $Z_{n}$.
@@ C.3 Semi-Synthetic Simulator
The second simulator is semi-synthetic, in the sense that the simulator parameters are learned based on real world data from the WSDM Cup, organised jointly by Spotify, WSDM, and CrowdAI [^4] https://research.atspotify.com/datasets/. The dataset comes from Spotify [^5] https://open.spotify.com/, which is an online music streaming platform with over 190 million active users interacting with a library of over 40 million tracks. The purpose of this challenge is to understand users’ behaviours based on sequential interactions with the streamed content they are presented with and how this contributes to skip track behaviour, which is considered as an important implicit feedback signal.
@@
For a detailed description of the dataset, please refer to Brost et al. [2019]. We created our simulator based on the “Test_Set.tar.gz (14G)” file. After unzipping the test data, we randomly chose a file named (“log_prehistory_20180918_000000000000.csv”) as our main data source. In our case, we only use the columns named: “session_id”, “session_position”, “session_length”, “track_id_clean”, and “skip_3”. Here, each unique “session_id” indicates a different sequence of interactions from an unknown user and the length of the interaction is denoted by “session_length”, where each step is denoted by “session_position”. The column “track_id_clean” indicates the particular track a user listen to at each position in the process. “skip_3” is a boolean value indicating if most of the track was played by the user. We assign the number of $1$ if it is “True” and 00 otherwise. We use this value to create $X_{n}^{t}$ later.
@@
To process the data, we start by looking at the existing sequence lengths and we select the most frequent sequence length in the data (20) (the actual corresponding session sequence is $10$), which counts for around $57.95\%$ of the original file. After filtering, there are around $228,460$ unique sessions and $301,783$ unique tracks in the data. To further reduce the computational cost of constructing a very large session-track matrix, we randomly sample a sub-population of $6000$ unique session ids, which then bring us to $26,707$ unique tracks. We build the session-track matrix where the entry value is the number of counts a particular track been listened over a particular session based on this sub-population. We apply singular-value decomposition (SVD) [^6] https://numpy.org/doc/stable/reference/generated/numpy.linalg.svd.html to this session-track matrix to get the session embeddings and track embeddings. For numerical stability reason, we save only the top $50$ singular values and further normalize it using its mean and standard deviation. For track embeddings, we take the top $10$ singular values after the normalization and assign them into $10$ clusters, using constrained $K$-means to to encourage equal-sized clusters [^7] https://github.com/joshlk/k-means-constrained. Using these clusters, we create a new categorical variable $F$ and its continuous counterparth $\phi(F)$, which respectively represent the corresponding cluster and its centroid vector based on the value of its associated “track_id_clean” entry. Next, we set $X_{n}^{t}$ as the cumulative sum of the corresponding binary “skip_3” value over each unique “session_id” with added zero-mean Gaussian noise with standard deviation $1/2$. We keep the first $5$ singular values for the session embedding normalization and consider them to be the session specific covariate $Z$. We save $X$, $F$, $\phi(F)$ and $Z$ as separate matrices as the final output, which have shapes of $6000\times 10$, $6000\times 10\times 10$, $6000\times 10\times 10$ and $6000\times 5$, respectively.
To build a simulator, we define the following two processes:
@@
Function $f(\cdot,\cdot,\cdot)$ is modeled with a deep neural network consisting of a single layer GRU and a MLP, trained on the real $F$ with cross entropy loss (as a likelihood based model). The second equation is trained on the real $X$ with an ordinary least square (OLS) regression with a ridge regularisation term ($\alpha=0.01$). Error variance is estimated by setting $\hat{\sigma}_{x}^{2}$ to be equal to the variance of the residuals. This defines a semi-synthetic model calibrated by our pre-processing of the real data.
We now need to define interventions synthetically, as they are not present on the real data. We choose $K=5$ as the size of the space of possible interventions. We will define interventions in a way to capture the notion of changing the exposure of tracks in particular cluster $k$ for particular user $n$ at time $t$. Since the real data does not contain information regarding how interventions can influence the behaviour of Spotify users, similar to the simulator setup for the fully-synthetic one, we use $5$ random seed to indicates different possible changes to the users behaviour in the Spotify platform.
@@
We set $w_{1d}=0$, $w_{2d}=0$ and $w_{3d}=1$ to reflect the “idle” or “default” intervention level $d=0$. The values are designed to symmetric along the default intervention value. We generate random time series $\{D_{n}^{t}\}$ based on the identification results in Section 2.2 and the rules we adopted in the full-synthetic simulator in Appendix C.2. Notice that the above does not correspond exactly to the model class we discuss in Section 2, as the product of intervention $\psi$ in Eq. 2 is multiplied to the softmax operation of $f(F_{n}^{1:t-1},X_{n}^{1:t-1},Z_{n})$ and also to the $\phi(F_{n}^{t})$.
We generate a synthetic dataset for training with the following parameters: $r=10$, $T=35$, $T_{0}=25$, $N=3000$, and $z_{\text{dim}}=5$ (note, we have the first $5000$ for training, but only use the first $3000$ samples are used for actual training, as we have additional experiments to check the influence of more training data points). For testing our performance, we generate additional non-overlapping $M=1000$ samples with the same parameters, but this time with $T=40$ for making predictions on the last $5$ time series based on the first $35$.
To sample data from this simulator, we take the initial state of $X_{n}^{0}$ and $F_{n}^{0}$ from the pre-processed data. We use the $Z_{n}$ vector which comes from the real-world user embedding. Then we iterative generated the next $F_{n}$ via categorical draw from the softmax output based on user embedding $Z_{n}$ all previous $F_{n}$ and $X_{n}$. Once a categorical value of $F_{n}$ is drawn, we replace it with its embedding form $\phi(F_{n})$ which is the centroid of its cluster. We use this $\phi(F_{n})$ to generate $X_{n}$ plus an additive Gaussian noise with zero mean and homoscedastic variance learned from data.
The outcome of this simulator gives us the following data: time series $\{X_{n}^{0:T}\}$, intervention sequence $\{D_{n}^{1:T}\}$ and covariates $Z_{n}$.

### v2

@@ Appendix B Proofs
$=Q_{C}(1-\alpha-{\rm gap}),$ (40)
where ${\rm gap}=\frac{\sqrt{N}}{Z}\sqrt{\sum_{n=1}^{N}\left(w_{n}^{t^{\prime}}-{\mathbb{E}}[w^{t^{\prime}}]\right)^{2}}=\frac{N}{Z}{\mathbb{V}}[w^{t^{\prime}}]=\frac{{\mathbb{V}}[w^{t^{\prime}}]}{{\mathbb{E}}[w^{t^{\prime}}]}$, with ${\mathbb{E}}[w^{t^{\prime}}]$ and ${\mathbb{V}}[w^{t^{\prime}}]$ being the empirical mean and variance of $w^{t}$ estimated on the $N$ calibration samples. Under the theorem assumptions, we have $w_{n}^{t^{\prime}}=(1-\epsilon)+\epsilon\frac{\delta(S_{n}^{t^{\prime}})}{p_{C}(S_{n}^{t^{\prime}})}$. Then, ${\mathbb{E}}[w^{t^{\prime}}]=1-\epsilon+\epsilon\frac{{\mathbb{E}}[\delta(S^{t^{\prime}})]}{p_{C}(S^{t^{\prime}})}\geq 1-\epsilon$ and ${\mathbb{V}}^{2}[w^{t^{\prime}}]={\mathbb{E}}[w^{t^{\prime}}-{\mathbb{E}}[w^{t^{\prime}}]]^{2}\leq\epsilon^{2}{\mathbb{V}}^{2}\left(\frac{\delta(S^{t^{\prime}})}{p_{C}(S^{t^{\prime}})}\right)\leq\frac{\epsilon^{2}}{(\min_{S}p_{C}(S^{t^{\prime}}))^{2}}{\mathbb{V}}^{2}[\delta(S^{t^{\prime}})]\leq\frac{\epsilon^{2}}{(\min_{S}p_{C}(S^{t^{\prime}}))^{2}}$. The statement follows from
${\rm Prob}(Z\leq a)\geq{\rm Prob}(Z\leq b)\quad{\rm if}\quad a\geq b,$ (41)
@@
$\alpha_{*}=\alpha+{\rm gap}\leq\alpha+\frac{\sqrt{{\mathbb{V}}^{2}[w^{t^{\prime}}]}}{{\mathbb{E}}[w^{t^{\prime}}]}\leq\frac{\epsilon}{\min_{S}p_{C}(S)}\frac{1}{1-\epsilon}.$ (42)
@@ Appendix C Simulators
We design two simulators: (1) a fully-synthetic simulator; and (2) a semi-synthetic simulator.
@@ C.1 General Setting
The simulators serve as fully controllable oracles to allow us to test the performance of our predictive causal inference problems. In particular, we have the following parameters:
@@
  • $r$: the dimensionality of the feature space.
  • $z_{\text{dim}}$: the number of time-invariant covariates.
@@
Whenever possible, we set the same random seeds of $1,2,3,4,5$ to aid reproducibility of results. For the fully-synthetic simulator, a different seed indicates that it is a different simulator (as given by random draws of simulator parameters based on this seed). This also reflects the randomness coming from neural network parameter initialization, and data splitting when training models. For the semi-synthetic simulator, the simulator parameters $\phi$ and $\beta$ are learned from real world data, but we will need to also draw synthetic parameters for $\psi$. In both cases, using different random seeds can be considered as drawing new problems where each has unique parameters.
@@ C.2 Fully-Synthetic Simulator
The first simulator is purely synthetic, where all parameters are randomly sampled from some pre-defined distribution. The overall data generation process exactly follows the model we specified in Section 2.
We generate a synthetic dataset for training with the following parameters: $r=5$, $T=20$, $T_{0}=10$, $K=5$, $N=50,000$, and $z_{\text{dim}}=5$. We generate $200,000$ samples at first place, but only use the first $50,000$ samples for training. To test our performance, we generate additional non-overlapping $M=5,000$ samples with the same parameters, but this time with $T=25$. We make predictions on the last $5$ time steps based on the first $20$.
@@
$w_{3d}\sim\mathcal{U}(2,3),$
where $\mathcal{U}(a,b)$ is the uniform distribution in $[a,b]$. For $d=3,4$:
@@
We set $w_{1d}=0$, $w_{2d}=0$ and $w_{3d}=1$ to reflect the “idle” or “default” intervention level $d=0$. Values are designed to symmetric around the default intervention value. We randomly generate time-series $\{D_{n}^{t}\}$ based on the identification results in Section 2.2, with the following rules:
@@
    for each unique unit, non-zero action levels are sampled without replacement. We additionally assume at least $I>K/2$ (in this case, $I=3$) of the interventions show up in each unit-level time series.
@@
    once a $d>0$ action happens, there must be at least two consecutive $d=0$ actions (see Theorem 1).
@@
    in the test dataset, at prediction time, we assume an unseen intervention in the first $20$ time series shows up at exactly step $21$, and then the time-series has intervention level of 0 until the end. For instance, the training sequence is $D=[0,0,\cdots,1,0,\cdots,0,2,0]$ with $T=20$, and test sequence is $D=[0,0,\cdots,4,0,\cdots,0,3,0]+[2,0,0,0,0]$ with $T=25$.
We generate a time series $\{\psi_{n}^{t}\}$, based on Eq. (2), using the corresponding $w_{1d},w_{2d},w_{3d}$ values. We generate $r$ iid entries $\beta_{nl}\sim\mathcal{N}(\mu_{n},\sigma_{n})$ for each user, as follows:
@@
For each user $n$, we generate $z_{\text{dim}}=5$ iid unit-specific covariates $\{Z_{n}\}$ as:
@@
The same is used to set up an initial $X_{n}^{0}$ (the initial starting point of the time series) with a random-parameter MLP with input dimension of $z_{\text{dim}}$ and output dimension of $1$. Finally, to create the synthetic time series $\{X_{n}^{0:T}\}$ (00 comes from the initial starting point), we randomly generate a neural network consisting of a single long-short term memory (LSTM) layer followed by a MLP to generate $\phi$ at each time step. We further compose $\phi$ with a sigmoid function so that it is bounded between $[-1,1]$. We iteratively sample the $\{X_{n}^{t}\}$ points with the conditional mean from Eq 1 (based on $X_{n}^{0:t-1}$, $Z_{n}$, and $D_{n}^{1:t}$) plus a standard Gaussian additive error term at each step.
The outcome of this simulator gives us the following data: time-series $\{X_{n}^{0:T}\}$, intervention sequence $\{D_{n}^{1:T}\}$ and covariates $Z_{n}$.
@@ C.3 Semi-Synthetic Simulator
The second simulator is semi-synthetic, in the sense that the simulator parameters are learned based on real world data from the WSDM Cup, organised jointly by Spotify, WSDM, and CrowdAI [^4] https://research.atspotify.com/datasets/. The dataset comes from Spotify [^5] https://open.spotify.com/, which is an online music streaming platform with over 190 million active users interacting with a library of over 40 million tracks. The purpose of this challenge is to understand users’ behaviours based on sequential interactions with the streamed content they are presented with, and how this contributes to skip track behaviour. The latter is considered as an important implicit feedback signal.
@@
For a detailed description of the dataset, please refer to Brost et al. [2019]. We created our simulator based on the “Test_Set.tar.gz (14G)” file. After unzipping the test data, we randomly chose a file named (“log_prehistory_20180918_000000000000.csv”) as our main data source. In our case, we only use the columns named: “session_id”, “session_position”, “session_length”, “track_id_clean”, and “skip_3”. Here, each unique “session_id” indicates a different sequence of interactions from an unknown user. The length of the interaction is denoted by “session_length”, where each step is denoted by “session_position”. The column “track_id_clean” indicates the particular track a user listens to at each position in the process. “skip_3” is a boolean value indicating whether most of the track was played by the user. We assign the number of $1$ if it is “True” and 00 otherwise. We use this value to later create $X_{n}^{t}$.
@@
To process the data, we start by looking at the existing sequence lengths, and select the most frequent sequence length in the data (20) (the actual corresponding session sequence is $10$). This accounts for approximately $58\%$ of the original file. After filtering, there are around $228,460$ unique sessions and $301,783$ unique tracks in the data. To further reduce the computational cost of constructing a very large session-track matrix, we randomly sample a sub-population of $6,000$ unique session ids, which then bring us to $26,707$ unique tracks. We build the session-track matrix where the entry value is the number of counts a particular track has been listened to over a particular session based on this sub-population. We apply singular-value decomposition (SVD) [^6] https://numpy.org/doc/stable/reference/generated/numpy.linalg.svd.html to this session-track matrix to get session embeddings and track embeddings. For numerical stability reasons, we save only the top $50$ singular values and further normalize them using its empirical mean and standard deviation. For track embeddings, we take the top $10$ singular values after the normalization and assign them into $10$ clusters, using constrained $K$-means to to encourage equal-sized clusters [^7] https://github.com/joshlk/k-means-constrained. Using these clusters, we create a new categorical variable $F$ and its continuous counterparth $\phi(F)$, which respectively represent the corresponding cluster and its centroid vector based on the value of its associated “track_id_clean” entry. Next, we set $X_{n}^{t}$ as the cumulative sum of the corresponding binary “skip_3” value over each unique “session_id” with added zero-mean Gaussian noise with standard deviation $1/2$. We keep the first $5$ singular values for the session embedding normalization and consider them to be the session specific covariate $Z$. We save $X$, $F$, $\phi(F)$ and $Z$ as separate matrices as the final output, which have shapes of $6,000\times 10$, $6,000\times 10\times 10$, $6,000\times 10\times 10$ and $6,000\times 5$, respectively.
To build the simulator, we further define the following two processes:
@@
Function $f(\cdot,\cdot,\cdot)$ is modeled with a deep neural network consisting of a single layer GRU and a MLP, trained on the real $F$ with cross entropy loss (as a likelihood based model). The second equation is trained on the real $X$ with an ordinary least square (OLS) regression with a ridge regularization term ($\alpha=0.01$). Error variance is estimated by setting $\hat{\sigma}_{x}^{2}$ to be equal to the variance of the residuals. This defines a semi-synthetic model calibrated by our pre-processing of the real data.
We now need to define interventions synthetically, as they are not present on the real data. We choose $K=5$ as the size of the space of possible interventions. We will define interventions in a way to capture the notion of changing the exposure of tracks in particular cluster $k$ for particular user $n$ at time $t$. Since the real data does not contain information regarding how interventions can influence the behaviour of Spotify users, similarly to the simulator setup for the fully-synthetic one, we use $5$ random seed to indicate different possible changes to a user behaviour in the Spotify platform.
@@
We set $w_{1d}=0$, $w_{2d}=0$ and $w_{3d}=1$ to reflect the “idle” or “default” intervention level $d=0$. The values are designed to symmetric around the default intervention value. We randomly generate time-series $\{D_{n}^{t}\}$ based on the identification results in Section 2.2 and the rules we adopted in the full-synthetic simulator in Appendix C.2. Notice that the above does not correspond exactly to the model class we discuss in Section 2, as the product of intervention features $\psi$ in Eq. (2) is pipeline with the softmax operation of $f(F_{n}^{1:t-1},X_{n}^{1:t-1},Z_{n})$ to get $F_{n}^{t}$, which is then composed with $\phi$ to get $\phi(F_{n}^{t})$.
We generate a synthetic training dataset with the following parameters: $r=10$, $T=35$, $T_{0}=25$, $N=3,000$, and $z_{\text{dim}}=5$ (notice that we have the first $5,000$ for training, but only use the first $3,000$ samples for the actual baseline training, as we have additional experiments to check the influence of more training data points). To test our performance, we generate additional non-overlapping $M=1,000$ samples with the same parameters, but this time with $T=40$ for making predictions on the last $5$ time-series steps based on the first $35$.
To sample data from this simulator, we take the initial state of $X_{n}^{0}$ and $F_{n}^{0}$ from the pre-processed data. We use the $Z_{n}$ vector which comes from the real-world user embedding. Then, we iteratively generate the next $F_{n}$ via categorical draws from the softmax output based on user embedding $Z_{n}$ all previous $F_{n}$ and $X_{n}$. Once a categorical value of $F_{n}$ is drawn, we replace it with its embedding form $\phi(F_{n})$, which is the centroid of its cluster. We use this $\phi(F_{n})$ to generate $X_{n}$ plus an additive Gaussian noise with zero mean and homoscedastic variance learned from data.
The outcome of this simulator gives us the following data: time-series $\{X_{n}^{0:T}\}$, intervention sequence $\{D_{n}^{1:T}\}$ and covariates $Z_{n}$.